OpenStudy (anonymous):

the number 84 is divisible by 2,3,4 and 6. true or false

OpenStudy (anonymous):

Nice video on the subject

OpenStudy (anonymous):

Find out. 84 / 2 = ? 84 / 3 = ? 84 / 4 = ? 84 / 6 = ? If you get a whole number, it is divisible, if you don't, it is not.